# What computer science has to say about the simulation hypothesis

### v4

@@ Appendix B The recursion theorem and Rice’s theorem
For any partial computable function $Q(x,y)$ there is a Turing machine with index $e$ such that $T^{e}(x)=Q(x,e)$ for all $x$ for which $Q(x,e)$ is defined.
@@
For any TM of the form $M(y,x)$ taking two arguments, define $\llbracket{M(y,x)}\rrbracket_{x}$ as the index $e$ such that $T^{e}(x)=M(y,x)$ for all $x$, with $y$ fixed to the indicated value.
@@
(Note that $S(.)$ is a total computable function.) Using this definition, choose an index $k$ such that
@@
for all $x,t$. (At least one such index must exist since the RHS of Eq. 44 is a partial computable function.)
@@
An important special case of the theorem is where $Q(.,.)$ is a total computable function. In this case the theorem says that that there must be an $e$ such that $T^{e}(x)=Q(x,e)$ for all $x$, and therefore $T$ is also total computable. So, we can restrict to total computable functions $Q(.,)$ and total TM’s $T^{e}$, as a special case. I call this extension the “total recursion theorem” in the main text.
Kleene’s second recursion theorem has elicited some interesting commentary. For example, [50] writes that
@@
Interestingly, Hartmanis didn’t encounter the recursion theorem until he was in his $20$’s — and yet went on to win the Turing award.
@@
So if there is any TM $T$ that computes a function that is not in $G$, then there must be a TM $T^{\prime}$ such that our algorithm cannot tell us whether the function that $T^{\prime}$ computes lies in $G$.
@@
An important special case is where $G$ is restricted to a set of binary-valued partial functions. Since each of those functions is represented as a bit string, the set comprises a language. In this case Rice’s theorem concerns the decidability of languages.
Appendix C Why $\Delta t$ is not a physical variable
@@
Appendix D Subtleties with the model in Section 2.6
Section 2.6 presented an example of how a portion of our actual universe could implement the purely formal model of universe $V$ given in Section 2.2. In that example the computer was implemented as a UTM, replete with tapes, etc. This might seem a particularly awkward way of modeling a portion of our cosmological universe whose dynamics is computationally universal. After all, our universe “runs in parallel”, whereas a UTM is a serial system. This means that to implement such a UTM would require copious use of energy barriers and the like, to prevent the parallel nature of Hamilton’s equations from “leaking through”.
@@
Appendix E Work in the philosophy of science literature on the simulation hypothesis
Much of philosophy of science, going back to Kant (at least), would be rendered moot if we are simulations. In particular, this would be the case for many of the flavors of “*** realism” (e.g., scientific realism), which suppose there “exists” some “real” physical truth which is not reducible to a mathematical formalism, a physical truth that instead is somehow “concrete”, and non-mathematical in its very essence — and so not just a computer algorithm. Indeed, the simulation hypothesis is consistent (in the formal logic sense) with all axiomatic systems for the laws of physics that have been offered. However, it true it would imply that there is no such “concrete”, non-mathematical reality. This establishes that it is impossible to establish the logical necessity of any of these flavors of realism, from all suggested axiomatic foundations to the laws of physics.
The central idea of the simulation hypothesis has been extended in an obvious manner, simply by noting that the aliens that simulate our universe might themselves be a simulation in the computer of some even more sophisticated species, and so on and so on, in a sequence of ever-more sophisticated aliens. Similarly, going the other direction, in the not too distant future it is conceivable that we would be able to produce our own simulation of a universe running in some future computer that we will create, a simulation complete with variables that constitute “conscious, reasoning agents”. Indeed, we might produce such a simulation in which the reasoning agents can produce their own simulated universe in turn, etc. If this were the case, then in the near future there might be a sequence of species’, each one with a computer running a simulation that produces the species just below it in the sequence, with us somewhere in the “middle” of that sequence.
The literature of the last few decades on the simulation hypothesis has focused almost exclusively on whether we, in our universe are such a simulation. This question is answered rather trivially if we adopt the view of ontic structural realism [45, 14], especially if that view is formalized in terms of Tegmark’s level IV multiverse [66, 67]: yes, in some universes we are a simulation, and no, in some other universes we are not. (See also [18, 55].)
These investigations of relative probabilities all beg (many) questions about what such probabilities might actually mean, formally. (Just as there are such questions concerning the idea of ascribing probabilities to the universes in a level IV multiverse in general.) How does one ascribe a measure, obeying the Kolmogorov axioms based on some associated sigma algebra, to a collection of events each of which is a universe? Note that it is not even clear that this collection would be a set, rather than a proper class of some sort. Directly reflecting such problems, one could not experimentally assess any proposed value of such a probability, e.g., with a proper scoring rule. [^27] One might try to use a Bayesian “degree of belief” interpretation of probability to circumvent this issue. However, without any decision that would be made based on the probability assigned, and associated loss function, one could not apply Bayesian decision theory. So it is hard to see how “degree of belief” is meaningful in this case. See [70] for some related arguments.
In particular, many papers considering such probabilities assume, implicitly or otherwise, that there is some way to assign a uniform probability to all universes. However, it is proven in Section 2.4 that this is mathematically impossible in sets with the cardinality of the kind of universes considered in these other papers. One might imagine that rather than a uniform probability distribution, one could assign some sort of Cantor (“fair coin”) measure to the set of possible universes, as is done for example in algorithmic information theory. However, it is proven in Section 7 that it is impossible to assign a Cantor measure to a set of universes obeying a natural set of restrictions.
Yet others have considered issues concerning the simulation hypothesis that might best be characterized as “moral” in nature [34, 23, 28, 40], This work considers potential normative implications if the simulation hypothesis holds, or even such implications that follow from the mere possibility that the simulation hypothesis holds.
Appendix F Semi-formal work that has been done on the CS theoretic aspects of the simulation hypothesis
The only work that has been done that is even tangentially related to the CS aspects of the simulation hypothesis was a set of semi-formal results presented in [73]. Some of those semi-formal results define “simulation” in terms of the relationship between languages at different levels of the time, polynomial, or exponential hierarchies of CS theory, and some of them define “simulation” in terms of the relationship between languages at different levels the arithmetic or analytic hierarchies of logic, or similar constructions. (See [24, 6, 32] for background on such hierarchies.) Other results in [73] define “simulation” in terms of the relationship between computational machines with different Turing degrees [65, 62]. Yet others consider the implications of Gödel’s second incompleteness theorem, defining simulation in terms of languages (in the logic theory sense) with nested sets of axioms.
None of these results in [73] consider what it means for a computational machine to simulate a physical universe, per se. There is no concern for “coupling” the mathematics of CS theory to the laws of physics of our universe. No work has been done to date concerning the simulation hypothesis that focuses on the dynamical system of an evolving physical universe, like ours.

### v5

@@ B.1 The recursion theorem
For any partial computable function $Q(x,y)$ there is a Turing machine with index $e$ such that $T^{e}(x)=Q(x,e)$ for all $x$.
@@
For any TM of the form $M(y,x)$ taking two arguments, define $\llbracket{M(y,x)}\rrbracket_{x}$ as the index $e$ such that $T^{e}(x)=M(y,x)$ for all $x$, with $y$ fixed.
Our goal is to find a way to satisfy this equation for $e=y$. In other words, we wish to find a $y$ so that
$T^{\llbracket{M(y,x)}\rrbracket_{x}}(x)=M(y,x)$ (37)
@@
Note that $S(.)$ is a total computable function. Using Eq. 38, choose one out of the infinite set of indices $k$ such that
@@
(Such an index must exist since the RHS of Eq. 39 is a partial computable function, and by running over $k$ the LHS runs over all TMs.)
@@
An important special case of the theorem is where $Q(.,.)$ is a total computable function. In this case the theorem says that that there must be an $e$ such that $T^{e}(x)=Q(x,e)$ for all $x$, and therefore $T^{e}(x)$ is also total computable. So as a special case of the recursion theorem we can restrict to total computable functions $Q(.,)$ and total TM’s $T^{e}$. I call this special case the “total recursion theorem” in the main text.
As an aside, the recursion theorem has elicited some pointed commentary. For example, [44] writes that
@@
Interestingly, Hartmanis didn’t encounter the recursion theorem until he was in his $20$’s — and yet despite the implications of his comment, he went on to win the Turing award.
B.2 Rice’s theorem
@@
So if there is any TM $T$ that (partially) computes a function that is not in $G$, then there must be a TM $T^{\prime}$ such that our algorithm cannot tell us whether the function that $T^{\prime}$ computes lies in $G$.
An important special case is where $G$ is restricted to a set of partial functions each of which outputs a single bit. In this case Rice’s theorem concerns the decidability of sets of languages.
Appendix C Appendix C: Why $\Delta t$ is not a physical variable
@@
Appendix D Appendix D: Subtleties with the model in Section 2.6
Section 2.6 presented an example of how a portion of our actual universe could implement the purely formal model of universe $V$ given in Section 2.2. In that example the computer was implemented as a UTM, replete with tapes, etc.. This might seem a particularly awkward way of modeling a portion of our cosmological universe whose dynamics is computationally universal. After all, our universe “runs in parallel”, whereas a UTM is a serial system. This means that to implement such a UTM would require copious use of energy barriers and the like, to prevent the parallel nature of Hamilton’s equations from “leaking through”.